solve.
(y^2-5y)^2-20(y^2-5y)-96=0

Answer by stanbon(75887)

solve.
(y^2-5y)^2-20(y^2-5y)-96=0
---------------------------------
[(y^2-5y)-24][(y^2-5y)-4] = 0
----
y^2-5y = 24 or y^2-5y=4
----
y^2-5y-24 = 0 or y^2-5y-4 = 0
(y-8)(y+3)=0 or y = [5+-sqrt(25+16)]/2
y = 8 or y = -3 or y = [5+-sqrt(41)]/2
======================================================
Cheers,
Stan H.